The sum of the first two terms of a geometric progression is 9 and the sum to infinity is 25. If the common ratio is positive, find the common ratio and the first term.   


1
Expert's answer
2022-05-04T12:07:57-0400
"a+ar=9"

"S=\dfrac{a}{1-r}=25, r>0"

"\dfrac{9}{1+r}=25(1-r)"

"25r^2=16"

"r_1=-\dfrac{4}{5}, r_2=\dfrac{4}{5}"

Since "r>0," we take "r=\dfrac{4}{5}."


"a=\dfrac{9}{1+\dfrac{4}{5}}=5"

"a=5, r=\dfrac{4}{5}"